# On the Reliability of Cue Conflict and Beyond

### 11 Mar 2026 v1

@@ A.10 Distorted Model Prediction in Cue-conflict
Appendix B Model Architecture Details
@@
Table B: Effect of model architecture on shape/texture sensitivity (with REFINED-BIAS dataset), bias (with cue-conflict dataset), and in-domain performance (i.e., top-1 accuracy on ImageNet-1k). SB and TB indicate relative shape and texture biases, respectively. The full and partial spaces denote full and partial decision spaces, respectively. The colors in the models column indicate the model family.
@@
Arch. Models In-domain Shape-Sens Texture-Sens SB TB SB TB
ViT ViT-S [11] 0.7860 0.7456 0.6671 0.2727 0.7273 0.2890 0.7110
ViT-B [11] 0.7898 0.7194 0.6418 0.3385 0.6615 0.3218 0.6782
Swin-S [34] 0.8290 0.7739 0.7505 0.2065 0.7935 0.2397 0.7603
Swin-B [34] 0.8307 0.8047 0.7506 0.2436 0.7564 0.2494 0.7506
CMT-S [22] 0.8331 0.8324 0.6085 0.3185 0.6815 0.2859 0.7141
CMT-B [22] 0.8452 0.8868 0.6228 0.3312 0.6688 0.2988 0.7012
CNN BagNet-9 [3] 0.4636 0.0517 0.7194 0.0498 0.9502 0.0650 0.9350
BagNet-17 [3] 0.5865 0.0995 0.8017 0.0076 0.9924 0.0484 0.9516
BagNet-33 [3] 0.6421 0.2386 0.8008 0.0186 0.9814 0.0622 0.9378
AlexNet [31] 0.5654 0.3617 0.5648 0.2000 0.8000 0.2636 0.7364
VGG-11 [46] 0.7037 0.4523 0.7358 0.0753 0.9247 0.1055 0.8945
VGG-13 [46] 0.7155 0.4337 0.7324 0.0820 0.9180 0.1152 0.8848
VGG-16 [46] 0.7337 0.5534 0.6672 0.0994 0.9006 0.1129 0.8871
VGG-19 [46] 0.7422 0.6130 0.6160 0.1373 0.8627 0.1532 0.8468
ResNet-18 [23] 0.6975 0.5741 0.5674 0.1913 0.8087 0.2367 0.7633
ResNet-34 [23] 0.7328 0.5680 0.5635 0.2438 0.7562 0.2595 0.7405
ResNet-50 [23] 0.7613 0.5630 0.5684 0.1944 0.8056 0.2221 0.7779
ResNet-101 [23] 0.7737 0.6590 0.5303 0.2700 0.7300 0.2866 0.7134
ResNet-152 [23] 0.7831 0.6378 0.5179 0.2612 0.7388 0.2940 0.7060
InceptionV3 [47] 0.7731 0.6944 0.5233 0.5583 0.4417 0.5038 0.4962
Appendix C Comparison of Model Training Strategies
@@ C.2 Confidence Intervals for Models with Different Training Strategies
Table D: Confidence intervals ($95\%$ CI) of texture bias and shape bias measured using REFINED-BIAS for models trained with different strategies. The colors in the Models column indicate the model family, consistent with the color scheme used in Fig. 6. The overlapping CIs indicate similar shape or texture bias, while non-overlapping CIs indicate the opposite.
@@ C.4 Visual Examples for Different Learning Strategies
Figure J: Examples of the RandAug method used in most mixed-augmentation models. The figure shows the effects of Color Jitter, Posterize, and Solarize augmentations employed by RandAug. The boxed regions in each image highlight the degree of texture change.
Figure K: Examples of texture degradation methods. The boxed regions in each image highlight the degree of texture change.
Figure L: Examples of augmentations used for each contrastive learning method. The parameters of all augmentations follow the original model settings. The boxed regions in each image highlight the degree of texture change.
Figure M: Examples of adversarial images generated by adversarial training methods under varying $l_{2}$ distance. The overall images are perceptually similar in both shape and texture compared to the original images.
Figure N: Examples of adversarial images generated by adversarial training methods under varying $l_{\infty}$ distance. The overall images are perceptually similar in both shape and texture compared to the original images.

### 12 Mar 2026 v2

@@ A.10 Distorted Model Prediction in Cue-conflict
A.11 Architecture Dependent Shape and Texture Trade-off
Figure J: Comparison of model architectures under a fixed learning strategy and their ImageNet-1k top-1 accuracy. (a) Model preferences are measured on the cue-conflict dataset using the preference metric. (b) Model shape sensitivities and (c) texture sensitivities are measured on the REFINED-BIAS dataset. The point size reflects the receptive field (RF) size of each model architecture, sourced from Araujo et al. [1]. $\rho$ denotes the Spearman’s correlation coefficient.
We observe an architecture-dependent shape–texture trade-off in the relationship between in-domain accuracy and cue sensitivity (Fig. 9c and d). We hypothesize that this behavior can be explained by differences in receptive field (RF) size across CNN architectures. In particular, we observe a positive correlation between RF size and shape sensitivity across CNN architectures (0.938 in Fig. Jb). This observation is consistent with prior analyses [11, 36, 62, 10, 40] suggesting that models with larger RF sizes capture broader spatial context, which may help integrate object-level structural information. Conversely, smaller RF size positively correlate with higher texture sensitivity (0.912 in Fig. Jc), which aligns with prior studies [12, 4] showing that models with smaller RF size tend to rely on local image features such as textures.
Taken together, these suggest that the RF size of CNNs can explain the observed shape–texture sensitivity trade-off. Although a similar trend appears in the cue-conflict benchmark (Fig. Ja), it arises only because the preference metric enforces the shape–texture trade-off by its design. Although these correlations do not strictly imply causality, they highlight the RF as a plausible underlying mechanism that partially explains the observed trade-off.
Appendix B Model Architectures
B.1 Details of Models Architectures
@@
Table B: Effect of training strategies on shape/texture sensitivity (with REFINED-BIAS dataset), bias (with cue-conflict dataset), and in-domain performance (i.e., top-1 accuracy on ImageNet-1k). SB and TB indicate relative shape and texture biases, respectively. RF denotes the receptive field size. The full and partial spaces denote full and partial decision spaces, respectively. The colors in the models column indicate the model family. The full and partial spaces denote full and partial decision spaces, respectively.
@@
Arch. Models RF In-domain Shape-Sens Texture-Sens SB TB SB TB
ViT ViT-S [13] - 0.7860 0.7456 0.6671 0.2727 0.7273 0.2890 0.7110
ViT-B [13] - 0.7898 0.7194 0.6418 0.3385 0.6615 0.3218 0.6782
Swin-S [37] - 0.8290 0.7739 0.7505 0.2065 0.7935 0.2397 0.7603
Swin-B [37] - 0.8307 0.8047 0.7506 0.2436 0.7564 0.2494 0.7506
CMT-S [24] - 0.8331 0.8324 0.6085 0.3185 0.6815 0.2859 0.7141
CMT-B [24] - 0.8452 0.8868 0.6228 0.3312 0.6688 0.2988 0.7012
CNN BagNet-9 [4] 9 0.4636 0.0517 0.7194 0.0498 0.9502 0.0650 0.9350
BagNet-17 [4] 17 0.5865 0.0995 0.8017 0.0076 0.9924 0.0484 0.9516
BagNet-33 [4] 33 0.6421 0.2386 0.8008 0.0186 0.9814 0.0622 0.9378
AlexNet [33] 167 0.5654 0.3617 0.5648 0.2000 0.8000 0.2636 0.7364
VGG-11 [50] 150 0.7037 0.4523 0.7358 0.0753 0.9247 0.1055 0.8945
VGG-13 [50] 156 0.7155 0.4337 0.7324 0.0820 0.9180 0.1152 0.8848
VGG-16 [50] 212 0.7337 0.5534 0.6672 0.0994 0.9006 0.1129 0.8871
VGG-19 [50] 268 0.7422 0.6130 0.6160 0.1373 0.8627 0.1532 0.8468
ResNet-18 [25] 435 0.6975 0.5741 0.5674 0.1913 0.8087 0.2367 0.7633
ResNet-34 [25] 899 0.7328 0.5680 0.5635 0.2438 0.7562 0.2595 0.7405
ResNet-50 [25] 427 0.7613 0.5630 0.5684 0.1944 0.8056 0.2221 0.7779
ResNet-101 [25] 971 0.7737 0.6590 0.5303 0.2700 0.7300 0.2866 0.7134
ResNet-152 [25] 1451 0.7831 0.6378 0.5179 0.2612 0.7388 0.2940 0.7060
InceptionV3 [51] 1311 0.7731 0.6944 0.5233 0.5583 0.4417 0.5038 0.4962
Appendix C Model Training Strategies
@@ C.2 Confidence Intervals for Models with Different Training Strategies
Table D: Confidence intervals ($95\%$ CI) of texture bias and shape bias measured using REFINED-BIAS for models trained with different strategies. The colors in the Models column indicate the model family. The overlapping CIs indicate similar shape or texture bias, while non-overlapping CIs indicate the opposite.
@@ C.4 Visual Examples for Different Learning Strategies
Figure K: Examples of the RandAug method used in most mixed-augmentation models. The figure shows the effects of Color Jitter, Posterize, and Solarize augmentations employed by RandAug. The boxed regions in each image highlight the degree of texture change.
Figure L: Examples of texture degradation methods. The boxed regions in each image highlight the degree of texture change.
Figure M: Examples of augmentations used for each contrastive learning method. The parameters of all augmentations follow the original model settings. The boxed regions in each image highlight the degree of texture change.
Figure N: Examples of adversarial images generated by adversarial training methods under varying $l_{2}$ distance. The overall images are perceptually similar in both shape and texture compared to the original images.
Figure O: Examples of adversarial images generated by adversarial training methods under varying $l_{\infty}$ distance. The overall images are perceptually similar in both shape and texture compared to the original images.